# The Latency Price of Threshold Cryptosystem in Blockchains

### v1

@@ 7.3 Evaluation Setup
Let $n$ denote the total stake before rounding, which is approximately $8.7\times 10^{8}$. The secrecy and reconstruction thresholds (in stakes) for the slow path are $t_{\sf sec}=0.5n$ and $t_{\sf rec}=0.660n$, respectively. The secrecy and reconstruction thresholds (in stakes) for the fast path are $t_{\sf sec}^{\prime}=0.667n$ and $t_{\sf rec}^{\prime}=0.830n$, respectively. The total weight of the mainnet validators after rounding is $244$. The weight threshold for the slow path is $w=143$, and that for the fast path is $w^{\prime}=184$.
@@
Evaluation metrics. We measure the randomness latency as the duration required to generate randomness for each block, as in Definition 12. It measures the duration from the moment the block is finalized by consensus to the when the randomness for that block becomes available. We report the average randomness latency (measured over a period of 12 hours). We also measure and compare the setup overhead for Das et al. [23] and fast-path, using microbenchmarks on machines of the same hardware specs as the Aptos mainnet. As mentioned in Section 7.2, fast-path requires the DKG to share the same secret in two different thresholds, resulting in approaximately twice the cost compared to Das et al. [23]. We also measure the end-to-end DKG latency of fast-path on the Aptos mainnet.
7.4 Evaluation Results.
We report the latency measurements observed on Aptos mainnet.
Randomness latency. Table 3 summarizes the latency comparison of our fast-path and Das et al. [23]. As observed, fast-path significantly reduces the randomness latency of Das et al., by 71%. As mentioned in Section 6.2, the small latency overhead of fast-path comes from the fact that honest parties may need to wait slightly longer after local finalization to receive additional shares from the fast path, since the reconstruction threshold of the fast path is higher than the finalization threshold.
Randomness latency
Das et al. [23] fast-path
85.5 24.7
Table 2: Latencies of Das et al. [23] and our fast-path on Aptos mainnet. The latency unit is millisecond.
@@
8 Discussions
As we discussed in §1, our results applies to many other threshold cryptosystems that are natively integrated into blockchains. Next, we provide a few specific examples.
Threshold signature for state certification. Modern blockchains [10, 47] typically require validators to sign the cryptographic digest of the blockchain state (e.g., the root of a Merkle tree representing the state) after executing transactions in each block to certify the state changes to external clients, and prevent forks caused by nondeterminism due to software bugs. Existing blockchains achieve this using either a (weighted) threshold signature or a multi-signature. However, this introduces an additional message delay in every transaction’s latency. For the (weighted) threshold signature or the multi-signature, our solution for blockchain-native threshold cryptosystem with tight thresholds (Algorithm 2) can be used to reduce this additional message delay as follows.
During the blockchain consensus protocol, validators start executing the transactions in a block immediately upon receiving the block, i.e., before the block is finalized. Once the block is prefinalized, each validator sends its partial signature on the blockchain state to all other validators. Algorithm 2 guarantees that any state certified by the aggregated signature will be finalized by the blockchain consensus protocol. We want to note that this approach may lead to wasted computation in executing blocks that do not get finalized by the consensus protocol (called orphan blocks). Nevertheless, in modern high-performance blockchains [10, 47], where orphan blocks are rare, we believe that this trade-off is justified.
@@ 8 Discussions
9 Conclusion
In this paper, we studied the problem of blockchain-native threshold cryptosystem and its associated latency overhead. Unlike prior works that introduces at least one additional message delay atop the underlying consensus protocol, we designed a simple protocol that eliminates this overhead for blockchain-native threshold cryptosystem with tight thresholds for a wide class of blockchain protocols. For blockchain-native threshold cryptosystem with ramp thresholds, we formally proved that the latency overhead incurred by prior works is unavoidable. To address this issue, we proposed a protocol that can significantly reduce this overhead in optimistic scenarios, such as when the network is synchronous and the adversary corrupts parties owning a very small fraction of the total shares. We demonstrated the performance improvement of our protocol by deploying it atop the distributed randomness scheme of the Aptos blockchain.
Acknowledgments

### v3

@@ 0.E.3 Evaluation Setup
Let $n$ denote the total stake before rounding, which is approximately $8.7\times 10^{8}$. The secrecy and reconstruction thresholds (in stakes) for the slow path are $t_{\sf sec}=0.5n$ and $t_{\sf rec}=0.66n$, respectively. The secrecy and reconstruction thresholds (in stakes) for the fast path are $t_{\sf sec}^{\prime}=0.67n$ and $t_{\sf rec}^{\prime}=0.83n$, respectively. The total weight of the mainnet validators after rounding is $244$. The weight threshold for the slow path is $w=143$, and that for the fast path is $w^{\prime}=184$.
@@
Evaluation metrics. We measure the randomness latency as the duration required to generate randomness for each block, as in Definition 9. It measures the duration from the moment the block is finalized by consensus to the when the randomness for that block becomes available. We report the average randomness latency (measured over a period of 12 hours). We also measure and compare the setup overhead for Das et al. [23] and fast-path, using microbenchmarks on machines of the same hardware specs as the Aptos mainnet. As mentioned in Section 0.E.2, fast-path requires the DKG to share the same secret in two different thresholds, resulting in approaximately twice the cost compared to Das et al. [23]. We also measure the end-to-end DKG latency of fast-path on the Aptos mainnet.
0.E.4 Evaluation Results.
@@
Appendix 0.F Discussions
As we discussed in Section 1, our results applies to many other threshold cryptosystems that are natively integrated into blockchains. Next, we provide one more specific example.